# Fast X-ray Transient detection with AXIS: application to Magnetar Giant Flares

### v1

@@ Abstract
Magnetar giant flares (MGFs) are among the most luminous high-energy transients in the local universe, consisting of a short, intense MeV $\gamma$-ray spike followed by a softer, pulsating X-ray tail and possibly delayed radioactive emission. While only three Galactic events have been firmly detected, several extragalactic candidates have recently been reported, motivating the need for sensitive, rapid-response $\gamma$- and X-ray facilities to constrain their rates and energetics. We present a feasibility study of detecting MGFs with the Advanced X-ray Imaging Satellite (AXIS), focusing on two complementary pathways: (i) serendipitous discovery of the prompt $\gamma$-ray spike within the field of view, and (ii) rapid follow-up of MGF tails in nearby galaxies. Using sensitivity rescaling and volumetric rate estimates, we find that serendipitous detection of prompt spikes during the mission lifetime is possible but unlikely, given their hardness and short duration. In contrast, AXIS’s superior sensitivity, if accompanied by fast repointing capabilities, offer an extraordinary opportunity to detect pulsating X-ray tails out to $\sim$20 Mpc, enabling the first extragalactic measurements of periodic modulations and potentially constraining emission geometry and fireball physics. Finally, we evaluate the detectability of soft X-ray line emission from r-process nucleosynthesis in MGFs, finding that such signals are extremely faint and confining the detection to Galactic distances. Our study offer a general framework for assessing the detectability of short transients with future missions.
@@ 1 Introduction
Among the possible engines powering fast X-ray transients (FXTs) are magnetars. These are the most highly magnetized neutron stars, with surface magnetic fields exceeding $10^{14}$ Gauss (Kouveliotou et al., 1998, 1999). Recent observational results and analysis suggest magnetar activity likely comprise at least some fraction of already-detected FXTs (e.g., Dillmann et al., 2025) or whose activity may be detectable as FXTs in the local universe (e.g., Hu et al., 2025; Pearlman et al., 2025).
Magnetars in our Galaxy typically reside in a quiescent state. From time to time they exhibit short-duration bursts (lasting milliseconds to seconds) and occasionally burst storms—a series of bursts occurring over minutes to days. These episodes often accompany a longer-lived outburst phase of higher persistent emission, which can last weeks to months (see Negro et al., 2024, for a recent review of magnetars transient activity).
Separately, giant flares (MGFs) represent rare, extreme, but not cataclysmic events, characterized by a brief, intense spike (lasting milliseconds) followed by a modulated tail that decays over several minutes. MGFs are likely associated with stellar crust activity (e.g., Liang & Antiochos, 1984; Thompson & Duncan, 1995; Lander et al., 2015; Lander, 2016, 2023; Bransgrove et al., 2025), possibly as the solid crust quakes caused by evolution and build-up of stresses by the strong magnetic field. Three nearby Galactic events have been observed (2 from the Milky Way, 1 from the LMC); while recently a total of 6 extragalactic MGFs candidates have been identified (Burns et al., 2021; Trigg et al., 2023, 2025). The spike reaches $E_{\rm iso}\sim 10^{44}-10^{46}$ erg and it’s energy spectrum, $E^{2}dN/dE$, peaks in the soft gamma rays, evolving in time from 1 MeV–0.3 MeV with a hard photon power-law index around $\sim$0).
@@
Figure 1 provides a schematic illustration the three phases of an MGF emission that we will consider in this work: the bright, short spike, the modulating decaying tail, and the radioactively powered (r-process) emission at later times.
@@
The Advanced X-ray Imaging Satellite (AXIS) (Reynolds et al., 2023) is a recently selected NASA Probe-class mission currently in Phase A. AXIS is designed to provide an order-of-magnitude improvement in sensitivity over Chandra, combined with a large field of view, rapid response, and excellent angular resolution. In this work we are interest in assessing AXIS capabilities in the context of FXTs, in particular, we consider all three phases of the MGFs emission and assess observational prospects.
@@
in Sec. 3, noting that, with a rapid response aĺa Swift, the superior sensitivity of AXIS could provide the first opportunity to probe MGFs’ periodic tail from local (<30<30 Mpc) extragalactic MGFs, we determine the optimal time-to-target (TtT) that would be required to detect and characterize such signals to extragalactic distances.
@@ 2 MGF serendipitous detection
AXIS is a pointed soft X-ray observatory, and its observing strategy is necessarily guided by predefined targets of interest on the sky. The unpredictable nature of transient phenomena—appearing at random times and locations—makes time-domain observations challenging. In practice, AXIS can observe such events under two circumstances: either the transient serendipitously occurs within the telescope’s field of view during a scheduled pointing, or, if the event duration is sufficiently long relative to the spacecraft’s slewing capabilities, AXIS can rapidly re-point in response to external alerts. For MGF spikes, however, the bright prompt emission typically lasts only a few milliseconds, leaving serendipitous coverage as the only viable detection channel. In this section, we therefore estimate the expected number of MGFs that may fall within the AXIS field of view by chance and assess the fraction that could be detectable as a function of the distance from us.
To do this we need some ingredients. First, Assuming a signal-dominated burst lasting 64 ms, we estimate the AXIS sensitivity for this transient detection to be $2.2\times 10^{-10}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$. This is obtained from the deep (7 Ms) field sensitivity in the 0.3–3 keV ($\sim 10^{-18}\,\mathrm{erg\,\,cm^{-2}\,s^{-1}}$, (Reynolds et al., 2023)) rescaled to 64 ms. Secondly, We assume the local ($z=0$) intrinsic volumetric rate as measured by Burns et al. (2021)  to be $3.8\times 10^{5}\,\,\mathrm{Gpc^{-3}\,yr^{-1}}$ for a spike of $E_{\rm iso}>3.7\times 10^{44}\,\mathrm{erg}$.
Because the MGF spectral and energetic quantities are bolometric, we need to know the conversion factor to go from bolometric quantities to AXIS energy range. Assuming a typical observed Comptonized spectrum of MGFs (power-law index of 0.6, peak energy 290 keV) the ratio between the integrated flux in AXIS 0.3-10 keV band and the bolometric (1-10000 keV) band is $f_{\rm bol2AXIS}\sim 0.014$. The maximum bolometric energy observed so far is $E_{iso,bol}\sim 5.75\times 10^{46}$ erg (Burns et al., 2021), corresponding to a maximum luminosity in the AXIS band (considering a duration of $\sim 64$ ms) of $L_{\rm iso,AXIS}^{\rm max}=L_{\rm iso,bol}^{\rm max}\times f_{\rm bol2AXIS}\sim 1.3\times 10^{46}$ erg/s, where $L_{\rm iso,bol}^{\rm max}=E_{\rm iso,bol}/64$ ms.
Considering effective observation factor of 95% (i.e., removing the time spent slewing/calibrating) and accounting for a 450 arcmin2 ($3.8\times 10^{-5}$ sr) field of view, we define an AXIS total observing efficiency as $\epsilon_{\rm AXIS}=0.95\Omega_{\rm FoV}$, where $\Omega_{\rm FoV}=\frac{\rm FoV[sr]}{4\pi}$.
@@
Figure 2: Left: Fraction of AXIX detectable events as function of the distance. Right: Cumulative number of MGFs as seen by AXIS in 5 years mission as a function of the distance.
@@
Multiplying by 5, corresponding to the 5-year prime mission duration, we can obtain the cumulative number of events out to the maximum distance AXIS would detect. Fig. 2 (right) shows the cumulative number of MGF events over 5 years that will fall within AXIS’s field of view (blue curve) and those that would be detected, being above AXIS the sensitivity threshold (orange curve). AXIS detection prediction in 5 years of mission is $N^{\rm MGFs}_{\rm AXIS\,\,5yr}=0.21^{+0.44}_{-0.04}$.
Note, this calculation ignores the greater hardness for brighter events. We assumed an average Comptonized spectrum (index = 0.6, energy peak = 250 keV), but a harder spectrum (i.e., index=0.0, energy peak = 1500 keV) can be expected for brigther events and it would result in a slightly lower $f_{bol2AXIS}$. In the worst case scenario that all MGFs have harder spectra (which we know is not the case) we get only a minor change in the final estimate of the detectable MGFs, namely $N^{\rm MGFs}_{\rm AXIS\,\,5yr}=0.19^{+0.39}_{-0.04}$. Also ignored in this calculation is the possible evolution of the rate with redshift. Such evolution is not known exactly, but assuming that MGFs trace massive star evolution, we estimate that accounting for this effect would only improve the serendipitous detection rate by a $\sim 10\%$.
To summarize, AXIS could detect all MGFs out to 55 Mpc, the 40% brightest ones out to 100 Mpc and the 10% brightest within 300 Mpc. Unfortunately, given the relatively local distances it is quite unlikely that AXIS could serendipitously observe such fast transients in its FoV in 5 years.
It is worth noting that AXIS would gain a significant advantage by devoting a fraction of its observing time to nearby galaxies (<<100 Mpc), particularly star-forming systems. In this regime, the volumetric average assumption breaks down, since a small number of galaxies with high star-formation activity can dominate the overall event rate (Beniamini et al., 2025). Targeted monitoring of such galaxies would therefore enhance the chances of capturing MGFs and, more generally, fast X-ray transients, serendipitously above the isotropic expectation. Within the Local Group, the Milky Way, LMC, M31, and M33 together contribute a combined star formation rate of about 3.5 $M_{\odot}$ yr-1. If gamma-ray detections and identifications are complete out to these distances—which is plausible, provided timely follow-up and robust identification strategies—the corresponding MGF rate would be approximately 0.2 events per year. This translates into the possibility of detecting roughly one MGF over a five-year mission: an exciting prospect, though still uncertain.
@@ 2.1 Generic framework to assess FXT detectability
Instrument needed information Transient needed information Our Assumptions
@@
Observing efficiency: combination of duty cycle and FoV coverage Luminosity/energy distribution (e.g., power-law) No cosmological evolution of event rate (unless modeled explicitly)
@@
The exercise made above for the case of MGFs could be applied to other types of short transients, where the term ‘short’ can be interpreted as broad as any transient with duration such that the assumption of negligible background still holds. AXIS’s background rate is estimated to be around $4.43\times 10^{-8}$ counts/s, which can be considered negligible ($10^{-3}$ counts) up to $10^{5}$ s exposure. Because this could be relevant to other groups, possibly beyond the specifics of AXIS we provide a summary of needed quantities in Tab. 1 and assumptions needed if one would like to repeat this calculation for different events and missions. The generic workflow looks like this:
@@
    Instrument characterization: Determine the sensitivity (flux or energy flux threshold), energy band, field of view, and effective observing factor (e.g., duty cycle, slewing losses).
@@
    Transient population properties: Specify the burst duration, intrinsic volumetric rate (local or redshift-dependent), maximum observed isotropic energy (or luminosity), and typical spectral shape with characteristic parameters (e.g., index, peak energy). Define the luminosity distribution of the event population (e.g., power-law).
@@
    Detection horizon: Compute the maximum detectable distance by comparing transient luminosity to the instrument sensitivity in the relevant energy band. Depending on the instrument bandpass this steps might need a bandpass correction from the bolometric quantities. This can be done considering the average spectrum of the population.
@@
    Detection fraction: For each distance, estimate the fraction of events above the sensitivity threshold, accounting for the luminosity distribution scaled to the distance.
@@ 3 MGF tails detection
Ref. Safi-Harb et al. (2023) estimated that, based on its sensitivity, AXIS could detect magnetar candidates out to a distance of $\sim$3.5 Mpc in a deep 250 ks exposure, via persistent emission. However, confirming such detections through periodicity searches would be extremely challenging in the absence of Athena (Barcons et al., 2012). At present, the most promising strategy appears to be the detection of the tail emission of a giant flare or an ‘intermediate flare’ outburst (Hu et al., 2025). The arcsecond-scale localization provided by AXIS would further enable studies of the environments of extragalactic magnetars, likely associated with star-forming regions, as observed in the Milky Way, but potentially with unexpected variations.
Given the high scientific impact resulting from being able to study extragalactic magnetars, we investigate the feasibility of repointing AXIS in response to externally triggered MGFs (i.e., detected by gamma-ray burst monitors). In particular, we assess how rapidly the observatory would need to slew in order to enable a robust detection [^1] That is, obtaining a signal clearly distinguishable from the background with a meaningful characterization. by collecting sufficient photons to identify the pulsations.
Figure 3: Left: Lightcurve of an MGF tail mimiking the one observed during the 2004 even from SGR 1806–20 (Hurley et al., 2005). Right: Simulated AXIS light curve of a decaying tail similar to the 2004 event from SGR 1806–20.
To determine the detectability of extragalactic MGFs tails by AXIS we take the observed tail of the 2004 giant flare from SGR 1806–20 (Hurley et al., 2005) and simulate an AXIS observation assuming the same spectrum and the one measured for the event and accounting for the AXIS energy dependent effective area.
The simulation assumes a modulated blackbody tail with a $5$ second period and an overall decay consistent with a trapped fireball scenario:
@@
The flux of the first peak was converted to AXIS count rate using webPIMMS [^2] https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3pimms/w3pimms.pl assuming an unabsorbed flux of $1.31\times 10^{-4}$ erg cm-2 s-1 (2010020-100 keV) as input [^3] The flux $1.31\times 10^{-4}$ erg cm-2 s-1 (20–100 keV) is derived from the intrinsic isotropic luminosity estimated at the beginning of the tail by Hurley et al. (2005) and assuming a distance of the magnetar of 8 kpc., a blackbody temperature of $kT=8$ keV, and a Galactic absorption column of $N_{H}=10^{20}$ cm-2, according to litterature values. webPIMMS estimates an initial count rate of $1.1\times 10^{6}$ counts/s on-axis (and $\sim 8.3\times 10^{5}$ counts/s averaged over the AXIS field of view) in the 0.3–10 keV band.
@@
The simulated AXIS-detected tail is then scaled to increasing extragalactic distances from 0.8 Mpc (M31) to 30 Mpc. For each case, Poisson fluctuations were added to the binned light curve. The resulting expected AXIS counts (per 10 s binning) are shown in Fig.4 (left) for representative galaxies: M31 (0.8 Mpc), NGC 253/M82/M81 ($\sim 3.7$ Mpc), M83 (4.5 Mpc), as well as for generic sources placed at 10, 20, and 30 Mpc. Also in this case the duration of the tail is short enough to allow us to assume a negligible background rate: $\sim 4\times 10^{-8}$ counts/s in a 0.6”-radius circular region around the transient location. The background was estimated from the AXIS simulated full-FoV background spectrum, integrated between 0.3–10 keV and rescaled for the selected region), a tail detection could happen up to $\sim 20$ Mpc with a fast repoint (á la Swift (Gehrels et al., 2004)). The power spectra are computed from the simulated AXIS event lists and adopting a 1 second binning. Fig. 4 (right) shows the power spectra for a periodic decaying tale from a MGF at different distances. The input signal at 5-seconds periodic modulation is evident, yielding a peak in the power spectra near 0.2 Hz. As distance increases and photon statistics degrade, the periodic signal becomes progressively less detectable. While the periodicity is prominent in nearby galaxies (e.g., M31 at 0.8 Mpc), it becomes indistinguishable from noise beyond $\sim 15-20$ Mpc.
@@
To consider is the time delay from the beginning of an MGF tail emission and the instant AXIS would be on target and start observing. We call such a delay time-to-target (TtT). In particular we are interested in the maximum TtT that would allow a detection of the typical second-log modulation of MGF tails as a function of the distance. Fig. 5 (right), shows exactly this. The orange curve indicates the maximum time within which a transient can be detected to confidently identify it as an MGF (minimum of $\sim$2 counts detected with zero background expected). The purple curve shows the more stringent condition required to detect periodic modulation in the tail at about 3 sigma (minimum of $\sim$40 counts detected). Vertical dashed lines mark reference distances of 10 Mpc and 20 Mpc. The step-like features and fluctuations in both curves arise from noise in the simulations.
We estimate that AXIS could identify an MGF tail out to $\sim 20$ Mpc detecting at least 2 cts, with a $TtT~\sim 200$ s, and can detect ($3\sigma$) modulation of MGF tail out to $\sim 5$ Mpc detecting at least 40 cts, with a $TtT\sim 150$ s. This fast turn-around is currently possible with the technology onboard the Neil Gehrels Swift Observatory (Gehrels et al., 2004).
Once again, we want to remark how adopting an observing strategy in which nearby galaxies (<<50 Mpc) fall in AXIS’ FoV, could boost the chance of detecting orphan tails serendipitously, offering crucial new observations to shed light on the physics, collimation/beaming, and geometry of these events.

### v2

@@ Abstract
Magnetar giant flares (MGFs) are among the most luminous high-energy transients in the local universe, consisting of a short, intense MeV $\gamma$-ray spike followed by a softer, pulsating X-ray tail and possibly delayed radioactive emission. While only three Galactic events have been firmly detected, several extragalactic candidates have recently been reported, motivating the need for sensitive, rapid-response $\gamma$- and X-ray facilities to constrain their rates and energetics. We present a feasibility study of detecting MGFs with the Advanced X-ray Imaging Satellite (AXIS), focusing on two complementary pathways: (i) serendipitous discovery of the prompt $\gamma$-ray spike within the field of view, and (ii) rapid follow-up of MGF tails in nearby galaxies. Using sensitivity rescaling and volumetric rate estimates, we find that serendipitous detection of prompt spikes during the mission lifetime is possible but unlikely, primarily because of their short duration and primarily because of their short duration and hard spectrum, in the assumption that the hard gamma-ray spectrum can be reliably extrapolated to the instrument’s energy range. In contrast, AXIS’s superior sensitivity, if accompanied by fast repointing capabilities, offer an extraordinary opportunity to detect pulsating X-ray tails out to $\sim$20 Mpc, enabling the first extragalactic measurements of periodic modulations from a magnetar and potentially constraining emission geometry and fireball physics. Finally, we evaluate the detectability of soft X-ray line emission from r-process nucleosynthesis in MGFs, finding that such signals are extremely faint and confining the detection to Galactic distances. Our study offer a general framework for assessing the detectability of short transients with future missions.
@@ 1 Introduction
Among the possible engines powering fast X-ray transients (FXTs) are magnetars. These are the most highly magnetized neutron stars, with surface magnetic fields inferred to exceed $10^{14}$ Gauss from timing measurements (Kouveliotou et al., 1998, 1999), consistent with theoretical expectations (Thompson & Duncan, 1995) and subsequent observational evidence across the X-ray and gamma-ray bands (see Mereghetti et al., 2015, and references therein). Recent observational results and analysis suggest magnetar activity likely comprise at least some fraction of already-detected FXTs (e.g., Dillmann et al., 2025) or whose activity may be detectable as FXTs in the local universe (e.g., Hu et al., 2025; Pearlman et al., 2025).
Magnetars in our Galaxy typically reside in a quiescent state. From time to time they exhibit short-duration bursts (lasting milliseconds to seconds) and occasionally burst storms—a series of bursts occurring over minutes to days. These episodes often accompany a longer-lived outburst phase of higher persistent emission, which can last weeks to months (see Negro et al., 2024, for a recent review of magnetars’ transient activity).
Separately, giant flares (MGFs) represent rare, extreme, but not cataclysmic events, characterized by a brief, intense spike (lasting milliseconds) followed by a modulated tail that decays over several minutes. MGFs are likely associated with stellar crust activity (e.g., Liang & Antiochos, 1984; Thompson & Duncan, 1995; Lander et al., 2015; Lander, 2016, 2023; Bransgrove et al., 2025), possibly as the solid crust quakes caused by evolution and build-up of stresses by the strong magnetic field. Three nearby Galactic events have been observed (2 from the Milky Way, 1 from the LMC); while recently a total of six extragalactic MGF candidates have been identified (Burns et al., 2021; Trigg et al., 2023, 2025; Roberts et al., 2021; Mereghetti et al., 2024; Rodi et al., 2025). The spike reaches $E_{\rm iso}\sim 10^{44}-10^{46}$ erg and its energy spectrum, $E^{2}dN/dE$, peaks in the soft gamma rays, evolving in time from 1 MeV to 0.3 MeV with a hard photon power-law index of $\sim$0).
@@
Figure 1 provides a schematic illustration of the three phases of an MGF emission that we will consider in this work: the bright, short spike, the modulating decaying tail, and the radioactively powered (r-process) emission at later times.
@@
The Advanced X-ray Imaging Satellite (AXIS) (Reynolds et al., 2023) is a recently selected NASA Probe-class mission currently in Phase A. AXIS is designed to provide an order-of-magnitude improvement in sensitivity over Chandra, combined with a large field of view, rapid response, and excellent angular resolution. In this work we are interested in assessing AXIS capabilities in the context of FXTs, in particular, we consider all three phases of the MGF emission and assess observational prospects.
@@
In Sec. 3, noting that, with a rapid response aĺa Swift, the superior sensitivity of AXIS could provide the first opportunity to probe MGFs’ periodic tail from local (<30<30 Mpc) extragalactic MGFs, we determine the optimal time-to-target (TtT) that would be required to detect and characterize such signals at extragalactic distances.
@@ 2 MGF serendipitous detection
Assuming a signal-dominated burst lasting 64 ms, we adopt the AXIS sensitivity for transient detection as $2.2\times 10^{-10}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$. This is obtained from the deep (7 Ms) field sensitivity in the 0.3–3 keV ($\sim 10^{-18}\,\mathrm{erg\,\,cm^{-2}\,s^{-1}}$, (Reynolds et al., 2023)) rescaled to 64 ms. We aim to estimate the number of MGFs AXIS could serendipitously detect within its FoV in the 5 years of prime mission, and the maximum distance at which they could be detected.
To do this, we assume the local ($z=0$) intrinsic volumetric rate as measured by Burns et al. (2021)  to be $3.8\times 10^{5}\,\,\mathrm{Gpc^{-3}\,yr^{-1}}$ for a spike of $E_{\rm iso}>3.7\times 10^{44}\,\mathrm{erg}$.
Firstly, because the MGF spectral and energetic quantities are bolometric, we need to know the conversion factor to go from bolometric quantities to AXIS energy range. Assuming a typical observed Comptonized spectrum of MGFs (power-law index of 0.6, peak energy 290 keV) the ratio between the integrated flux in AXIS 0.3-10 keV band and the bolometric (1-10000 keV) band is $f_{\rm bol2AXIS}\sim 0.014$. The maximum bolometric energy observed so far is $E_{iso,bol}\sim 5.75\times 10^{46}$ erg (Burns et al., 2021), corresponding to a maximum luminosity in the AXIS band (considering a duration of $\sim 64$ ms) of $L_{\rm iso,AXIS}^{\rm max}=L_{\rm iso,bol}^{\rm max}\times f_{\rm bol2AXIS}\sim 1.3\times 10^{46}$ erg/s, where $L_{\rm iso,bol}^{\rm max}=E_{\rm iso,bol}/64$ ms. We note that the assumption that the spectrum in the soft X-ray band is the extrapolation of the gamma-ray Comptonized spectrum is an educated guess, but no direct observations are available. As pointed out in Rea et al. (2013), observations of ionospheric disturbances from the SGR 1900+14 giant flare suggest that the soft X-ray flux (3–10 keV) can exceed the hard X-ray level by nearly an order of magnitude (Inan et al., 1999), thus our assumption can be considered conservative.
Secondly, considering an effective observation factor of 95% (i.e., removing the time spent slewing/calibrating) and accounting for a 450 arcmin2 ($3.8\times 10^{-5}$ sr) field of view, we define an AXIS total observing efficiency as $\epsilon_{\rm AXIS}=0.95\Omega_{\rm FoV}$, where $\Omega_{\rm FoV}=\frac{\rm FoV[sr]}{4\pi}$.
@@
Figure 2: Left: Fraction of AXIS detectable events as function of the distance. Right: Cumulative number of MGFs as seen by AXIS in 5 years mission as a function of the distance.
@@
Multiplying by 5, corresponding to the 5-year prime mission duration, we can obtain the cumulative number of events out to the maximum distance AXIS would detect. Fig. 2 (right) shows the cumulative number of MGF events over 5 years that will fall within AXIS’s field of view (blue curve) and those that would be detected, being above the AXIS sensitivity threshold (orange curve). AXIS detection prediction in 5-years of mission is $N^{\rm MGFs}_{\rm AXIS\,\,5yr}=0.21^{+0.44}_{-0.04}$.
Note, this calculation ignores the greater hardness for brighter events. We assumed an average Comptonized spectrum (index = 0.6, energy peak = 250 keV), but a harder spectrum (i.e., index=0.0, energy peak = 1500 keV) would result in a slightly lower $f_{bol2AXIS}$. In the worst-case scenario that all MGFs have harder spectra (which we know is not the case) we get only a minor change in the final estimate: $N^{\rm MGFs}_{\rm AXIS\,\,5yr}=0.19^{+0.39}_{-0.04}$. Also ignored in this calculation is the possible evolution of the rate with redshift. Such evolution is not known exactly, but assuming that MGFs trace massive star evolution, we estimate that accounting for this effect would only improve the serendipitous detection rate by a $\sim 10\%$.
In conclusion, AXIS could detect all MGFs out to 55 Mpc, the brightest 40% out to 100 Mpc and the brightest 10% within 300 Mpc. Unfortunately, given the relatively local distances it is quite unlikely that AXIS could serendipitously observe such fast transients in its FoV in 5 years.
It is worth noting that AXIS would gain a significant advantage by devoting a fraction of its observing time to nearby galaxies (<<100 Mpc), particularly star-forming systems. In this regime, the volumetric average assumption breaks down, since a small number of galaxies with high star-formation activity can dominate the overall event rate (Beniamini et al., 2025). Targeted monitoring of such galaxies would therefore enhance the chances of capturing MGFs and, more generally, fast X-ray transients, serendipitously above the isotropic expectation. Within the Local Group, the Milky Way, LMC, SMC, M31, and M33 together contribute a combined star formation rate of about 3.5 $M_{\odot}$ yr-1. If gamma-ray detections and identifications are complete out to these distances—which is plausible, provided timely follow-up and robust identification strategies—the corresponding MGF rate would be approximately 0.2 events per year. This translates into the possibility of detecting roughly one MGF over a five-year mission: an exciting prospect, though still uncertain.
@@ 2.1 Generic framework to assess FXT detectability
Instrument information needed Transient information needed Assumptions
@@
Observing efficiency: combination of duty cycle and FoV coverage Luminosity/energy distribution (e.g., power-law index) No cosmological evolution of event rate (unless modeled explicitly)
@@
The exercise made above for the case of MGFs could be applied to other type of short transients, where the term ‘short’ can be interpreted as broadly as any transient with duration such that the assumption of negligible background still holds. AXIS’s background rate is estimated to be around $4.43\times 10^{-8}$ counts/s, which can be considered negligible ($10^{-3}$ counts) up to $10^{5}$ s exposure. Because this could be relevant to other groups, possibly beyond the specifics of AXIS we provide a summary of needed quantities in Tab. 1 and assumptions needed if one would like to repeat this calculation for different events and missions. The generic workflow looks like this:
@@
    Instrument characterization: Determine the sensitivity (flux threshold), energy band, field of view, and effective observing factor (e.g., duty cycle, slewing losses). Combine these into an overall observing efficiency.
@@
    Transient population properties: Specify the burst duration, intrinsic volumetric rate (local or redshift-dependent), maximum observed isotropic energy (or luminosity), and typical spectral shape with characteristic parameters (e.g., index, peak energy). Define the luminosity or energy distribution (e.g., power-law).
@@
    Detection horizon: Compute the maximum detectable distance by comparing transient luminosity to the instrument sensitivity in the relevant energy band.
@@
    Detection fraction: For each distance, estimate the fraction of events above the sensitivity threshold, accounting for the luminosity/energy distribution.
@@ 3 MGF tails detection
Ref. Safi-Harb et al. (2023) estimated that, based on its sensitivity, AXIS could detect magnetar candidates out to a distance of $\sim$3.5 Mpc in a deep 250 ks exposure, via persistent emission. However, confirming such detections through periodicity searches would be extremely challenging in the absence of Athena (Cruise et al., 2025). At present, the most promising strategy appears to be the detection of the tail emission of a giant flare or an ‘intermediate flare’ outburst (Hu et al., 2025). The arcsecond-scale localization provided by AXIS would further enable studies of the environments of extragalactic magnetars, likely associated with star-forming regions, as observed in the Milky Way, but potentially with unexpected variations.
Given the high scientific impact resulting from being able to study extragalactic magnetars, we investigate the feasibility of repointing AXIS in response to externally triggered MGFs detected by gamma-ray burst monitors. In particular, we assess how rapidly the observatory would need to slew in order to enable both a robust detection [^1] That is, obtaining a signal clearly distinguishable from the background with a meaningful characterization. by collecting sufficient photons to identify the pulsations.
Figure 3: Left: Lightcurve of an MGF tail mimicking the one observed during the 2004 event from SGR 1806–20 (Hurley et al., 2005). Right: Simulated AXIS light curve of a decaying tail similar to the 2004 event from SGR 1806–20.
To determine the detectability of extragalactic MGF tails by AXIS we take the observed tail of the 2004 giant flare from SGR 1806–20 (Hurley et al., 2005) and simulate an AXIS observation assuming the same spectrum and the one measured for the event and accounting for the AXIS energy dependent effective area.
The simulation assumes a modulated blackbody tail with a $5$ second period, a pulse fraction of 100% [^2] In the literature (Hurley et al., 2005; Frederiks et al., 2007) the tail is described as “highly modulated”, but no quantitative estimate of the pulse fraction is reported., and an overall decay consistent with a trapped fireball scenario:
@@
The flux of the first peak was converted to AXIS count rate using webPIMMS [^3] https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3pimms/w3pimms.pl assuming an unabsorbed flux of $1.31\times 10^{-4}$ erg cm-2 s-1 (2010020-100 keV) as input [^4] The flux $1.31\times 10^{-4}$ erg cm-2 s-1 (20–100 keV) is derived from the intrinsic isotropic luminosity estimated at the beginning of the tail by Hurley et al. (2005) and assuming a distance of the magnetar of 8 kpc. We note that this distance is debated in literature. We adopted the value 8 kpc as estimated in Bibby et al. (2008), which is nearer than other values found in previous literature. Because assuming a closer source results in a lower intrinsic luminosity, our choice can be considered conservative., a blackbody temperature of $kT=8$ keV, and a Galactic absorption column of $N_{H}=10^{20}$ cm-2, according to literature values. webPIMMS estimates an initial count rate of $1.1\times 10^{6}$ counts/s on-axis (and $\sim 8.3\times 10^{5}$ counts/s averaged over the AXIS field of view) in the 0.3–10 keV band.
@@
The simulated AXIS-detected tail is then scaled to increasing extragalactic distances from 0.8 Mpc (M31) to 30 Mpc. For each case, Poisson fluctuations were added to the binned light curve. The resulting expected AXIS counts (per 10 s binning) are shown in Fig.4 (left) for representative galaxies: M31 (0.8 Mpc), NGC 253/M82/M81 ($\sim 3.7$ Mpc), M83 (4.5 Mpc), as well as for generic sources placed at 10, 20, and 30 Mpc. Also in this case the duration of the tail is short enough to allow us to assume a negligible background rate: $\sim 4\times 10^{-8}$ counts/s in a 0.6”-radius circular region around the transient location. The background was estimated from the AXIS simulated full-FoV background spectrum, integrated between 0.3–10 keV and rescaled for the selected region). With a fast repointing similar to the Neil Gehrels Swift Observatory (Gehrels et al., 2004), tail detections could be possible up to $\sim 20$ Mpc. The power spectra are computed from the simulated AXIS event lists and adopting a 0.2 second binning, assuming AXIS time resolution to be $\sim$5 frames/s (Reynolds et al., 2023). Fig. 4 (right) shows the power spectra for a periodic decaying tail from a MGF at different distances. The input signal at 5-second periodic modulation is evident, yielding a peak in the power spectra near 0.2 Hz. As distance increases and photon statistics degrade, the periodic signal becomes progressively less detectable. While the periodicity is prominent in nearby galaxies (e.g., M31 at 0.8 Mpc), it becomes indistinguishable from noise beyond $\sim 15-20$ Mpc.
@@
To consider is the time delay from the beginning of an MGF tail emission and the instant AXIS would be on target and start observing. We call this delay the time-to-target (TtT). In particular we are interested in the maximum TtT that would allows a detection of the typical seconds-long modulation of MGF tails as a function of the distance. Fig. 5 (right), shows exactly this. The orange curve indicates the maximum time within which a transient can be detected to confidently identify it as an MGF (minimum of $\sim$2 counts detected with zero background expected). The purple curve shows the more stringent condition required to detect periodic modulation in the tail at about 3 sigma (minimum of $\sim$40 counts detected). Vertical dashed lines mark reference distances of 10 Mpc and 20 Mpc. The step-like features and fluctuations in both curves arise from noise in the simulations.
AXIS can identify an MGF tail out to $\sim 20$ Mpc detecting at least 2 counts, with a $TtT\sim 200$ s, and can detect ($3\sigma$) modulation of MGF tail out to $\sim 5$ Mpc detecting at least 40 cts, with a $TtT\sim 150$ s. Such fast turnaround is currently possible with the technology onboard the Neil Gehrels Swift Observatory.
Once again, we want to remark how adopting an observing strategy in which nearby galaxies (<<50 Mpc) fall in AXIS’s FoV, could boost the chance of detecting orphan tails serendipitously, offering crucial new observations to shed light on the physics, collimation/beaming, and geometry of these events.